# Analysis and Optimization of RIS-Assisted Cell-Free Massive MIMO NOMA Systems

### v1

@@ III Achievable Downlink Spectral Efficiency
$\phantom{x}\hskip 21.0pt\text{I}^{R}_{kn^{\prime}}=\left|\sum\limits_{m=1}^{M}\sqrt{\eta_{mkn^{\prime}}\rho_{d}}\left({u}_{mkn}-{\hat{u}}_{mkn}\right){\hat{z}}_{mk}^{*}\right|^{2}$,
$\phantom{x}\hskip 21.0pt\text{I}^{C}_{k^{\prime}n^{\prime}}=\sum\limits_{m=1}^{M}\left|\sqrt{\eta_{mk^{\prime}n^{\prime}}\rho_{d}}{u}_{mkn}{\hat{z}}_{mk^{\prime}}^{*}\right|^{2}$.
@@
After substituting (A)-(41) into (16), the closed-from SINR $\Gamma_{kn}$ can be expressed as (18) (shown at the top of page $6$).
@@ III-A1 Effect of imperfect SIC
Unlike previous works in cell-free mMIMO NOMA [20],[24] which considered perfect SIC, the presence of imperfect SIC introduces residual intra-cluster interference, denoted as $\text{IaCI}^{R}_{kn^{\prime}}$ (fourth term in the denominator of $\Gamma_{kn}$ in (18)). Moreover, the users can not perform perfect SIC in a single cluster due to the unavailability of instantaneous downlink CSI, intra-cluster pilot contamination, and channel estimation errors. Recall the residual intra-cluster interference from (A)
@@ IV-A Successive QT-based Power Allocation Optimization
1:Initialize $\{\boldsymbol{\xi}_{m}\}_{m=1}^{M}$ with equal power allocation (EPA), and set a maximum number of inner and outer iterations as $J_{2}$ and $J_{1}$, where ${\zeta}$ is a pre-defined threshold that decides the error tolerance between two adjacent iterations.
2:for $j_{1}$ = $1,2,\cdots,J_{1}$ do
@@
4:         for $j_{2}$ = $1,2,\cdots,J_{2}$ do
5:             For a given $\boldsymbol{\xi}_{m}^{j_{2}}$, compute the auxiliary variable $w_{kn}^{*}$ using (33).
@@
8:             Repeat until convergence $\lVert{\boldsymbol{\xi}_{m}^{(j_{1},j_{2})}-\boldsymbol{\xi}_{m}^{(j_{1},j_{2}-1)}\rVert}^{2}\leq{\zeta}$.
@@
11:    Repeat until convergence $\lVert{\boldsymbol{\xi}^{(j_{1},j_{2})}-\boldsymbol{\xi}^{(j_{1}-1,j_{2})}\rVert}^{2}\leq{\zeta}$.
@@ V Simulation Results
In this section, we evaluate the performance of the RIS-assisted cell-free mMIMO NOMA system using extensive simulations. For simulation, we consider a large geographical area of $(1000\times 1000)$ m2, where the APs and users are uniformly distributed. The four vertices of the considered region in terms of $(x,y)$ coordinates are $[-500,500]$ m, $[-500,-500]$ m, $[500,-500]$ m and $[500,500]$ m. The RIS is located at $(x,y)=(0,100)$ m in the considered simulation area. The users are deployed outside a semi-circle of radius $d_{\text{min}}=100$ m from the RIS. We consider the large-scale path-loss parameters according to the three-slope propagation model as $\beta_{mk}=\text{PL}_{mk}+\sigma_{sh}z_{mk}$ (dB[1], where $\text{PL}_{mk}$ is the distance dependent path-loss which depends on the distance between the transmitter and the receiver, $\sigma_{sh}z_{mk}$ represents the shadow fading with standard deviation $\sigma_{sh}$, and $z_{mk}\sim\mathcal{CN}(0,1)$. The shadow fading has a log-normal distribution with a standard deviation ($\sigma_{sh}$) equal to $14$ dB and $8$ dB for the direct and indirect links, respectively. The distance thresholds for the three slopes are $10$ m and $50$ m. The height of the APs, RIS, and users is $15$ m, $30$ m, and $1.65$ m, respectively [13].
@@ V-A Validation of closed-form spectral efficiency expression derived in Theorem-1
Fig. 2(a) illustrates the SE performance for different transmit powers $\rho_{d}$ while serving a total $20$ number of users with two different setups, i.e., $K=10$ clusters with $N=2$ users per cluster and $K=5$ clusters with $N=4$ users per cluster. We set $M=64$ APs, $L=64$ RIS elements, and the coherence interval, $\tau_{c}=100$. We observe from Fig. 2(a) that the derived closed-form SE expression using (18) (labeled as “Closed-form SE (Analytical)”) exactly matches with the SE obtained using (16) (labeled as “Closed-form SE (Simulation)”), which confirms the accuracy of the derivation. We also observe that the derived SE expression using (18) in closely follows the ergodic SE in (13) (labeled as “Ergodic SE (Simulation)”), which validates Theorem 1. To show the advantages of using RIS, we also plot the SE in Fig. 2(a) in the absence of RIS (labeled as “Closed-from SE without RIS”). We clearly observe that for transmitting power $\rho_{d}<0$ dBm, the SE without RIS is almost zero due to the high path loss and shadowing. On the contrary, with the integration of RIS, we get a significant boost in the SE value even with very low transmit power. Therefore, this justifies using the RIS when there is a weak direct link between the user and the AP. We also observe that for higher transmit power, i.e., $\rho_{d}\geq 40$ dBm, the SE obtained with and without RIS are equal. This is because at higher transmit powers, the direct link becomes dominant over the RIS-aided link. As we increase $N$, the SINR at each user decreases due to the increase in intra-cluster interference, which in turn reduces the SE.
V-B Impact of the number of users on the spectral efficiency
@@ V-C Comparison of cell-free NOMA and cell-free OMA
V-D Joint impact of the number of APs and RIS elements on the spectral efficiency
@@ V-F NOMA-OMA comparison with the proposed algorithm
Figure 5: a) Spectral efficiency versus AP transmit power for different direct path scenarios, b) Spectral efficiency versus AP transmit power for different user distribution scenarios, c) Convergence analysis of Algorithm $3$.
V-G Impact of the number of RIS elements on the spectral efficiency with phase optimization
@@ V-H Impact of shadow fading coefficient $\sigma_{sh}$ and $d_{\text{min}}$
V-I Convergence analysis
To analyze the convergence of the proposed algorithm we now plot in Fig. 5(c) the SE versus the number of iterations required for the Algorithm $3$ to converge, by varying the AP transmit power. For this study, we set $M=64$ APs, $K=10$ clusters, $N=2$ users in each cluster and $L=64$ RIS elements. We observe from Fig. 5(c) that Algorithm $3$ requires a small number of iterations (i.e., around $18$ iterations) to converge.
@@ V-K Comparison of RIS assisted cell-free NOMA with SDMA and TDMA
V-L SE versus number of users for different coherence interval
@@ VI Conclusion
$=\beta_{mkn}+\text{Tr}\left(\mathbf{\Theta}_{r}\mathbb{E}\left\{\mathbf{g}_{mr}\mathbf{g}_{mr}^{H}\right\}\mathbf{\Theta}_{r}^{H}\mathbb{E}\left\{\mathbf{h}_{rkn}\mathbf{h}_{rkn}^{H}\right\}\right),$
@@
$\text{IaCI}^{R}_{kn^{\prime}}$
$=\rho_{d}\!\sum_{n^{\prime}=n+1}^{N}\!\sum\limits_{m=1}^{M}\!\!\eta_{mkn^{\prime}}\!\!\left(\mathbb{E}\left\{\left|({u}_{mkn}{\hat{z}}_{mk}^{*}\right|^{2}\right\}-\left|\mathbb{E}\left\{{u}_{mkn}{\hat{z}}_{mk}^{*}\right\}\right|^{2}\right)\!$
=$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{n^{\prime}=n+1}^{N}\sum\limits_{m=1}^{M}{{\eta_{mkn^{\prime}}\rho_{d}}}\delta_{mkn}\gamma_{mk},$ (40)
@@
$\text{ICI}_{k^{\prime}n^{\prime}}=\sum_{k^{\prime}\neq k}^{K}\sum_{n^{\prime}=1}^{N}\sum\limits_{m=1}^{M}{\eta_{mk^{\prime}n^{\prime}}\rho_{d}}\underbrace{\mathbb{E}\left\{\left|{u}_{mkn}{\hat{z}}_{mk^{\prime}}^{*}\right|^{2}\right\}}_{Q},$
[…]

### v2

@@ III Achievable Downlink Spectral Efficiency
$\phantom{x}\text{I}^{R}_{kn^{\prime}}={\rho_{d}}\left|\sum\limits_{m=1}^{M}\!\!\!\sqrt{\eta_{mkn^{\prime}}}u_{mkn}\hat{z}_{mk}^{*}\!\!-\mathbb{E}\!\left\{\!\sum\limits_{m=1}^{M}\!\!\sqrt{\eta_{mkn^{\prime}}}u_{mkn}\hat{z}_{mk}^{*}\!\!\right\}\!\right|^{2}\!$,
$\phantom{x}\text{I}^{C}_{k^{\prime}n^{\prime}}=\rho_{d}\left|\sum\limits_{m=1}^{M}\sqrt{\eta_{mk^{\prime}n^{\prime}}}{u}_{mkn}{\hat{z}}_{mk^{\prime}}^{*}\right|^{2}$.
@@
After substituting (A)-(41) into (16), the closed-form SINR $\Gamma_{kn}$ can be expressed as (18) (shown at the top of page $6$).
@@ III-A1 Effect of imperfect SIC
Unlike previous works in cell-free mMIMO NOMA [20],[24] which considered perfect SIC, the presence of imperfect SIC introduces residual intra-cluster interference, denoted as $\text{IaCI}^{R}_{kn^{\prime}}$ (fourth term in the denominator of $\Gamma_{kn}$ in (18)). Moreover, the users can not perform perfect SIC in a single cluster due to the unavailability of instantaneous downlink CSI, intra-cluster pilot contamination, and channel estimation errors. Recall the residual intra-cluster interference from (40)
@@ IV-A Successive QT-based Power Allocation Optimization
1:Initialize $\{\boldsymbol{\xi}_{m}\}_{m=1}^{M}$ with equal power allocation (EPA), and set a maximum number of inner and outer iterations as $J_{i}$ and $J_{o}$, where ${\zeta}$ is a pre-defined threshold that decides the error tolerance between two adjacent iterations.
2:for $j_{o}$ = $1,2,\cdots,J_{o}$ do
@@
4:         for $j_{i}$ = $1,2,\cdots,J_{i}$ do
5:             For a given $\boldsymbol{\xi}_{m}^{j_{i}}$, compute the auxiliary variable $w_{kn}^{*}$ using (33).
@@
8:             Repeat until convergence $\lVert{\boldsymbol{\xi}_{m}^{(j_{o},j_{i})}-\boldsymbol{\xi}_{m}^{(j_{o},j_{i}-1)}\rVert}^{2}\leq{\zeta}$.
@@
11:    Repeat until convergence $\lVert{\boldsymbol{\xi}^{(j_{o},j_{i})}-\boldsymbol{\xi}^{(j_{o}-1,j_{i})}\rVert}^{2}\leq{\zeta}$.
@@ IV-C Computational Complexity
IV-D Convergence Analysis
We now analyze the convergence of the proposed algorithm using the following lemma.
Lemma 1.
Consider the optimization problems $\mathbf{P1A}$, $\mathbf{P2A}$, $\mathbf{P3A}$, and $\mathbf{P4A}$ with corresponding objective functions $f_{1A}(\boldsymbol{\eta})$, $f_{2A}(\boldsymbol{\eta})$, $f_{3A}(\boldsymbol{\eta}_{m},\boldsymbol{\eta}_{m}^{\mathsf{c}})$, and $f_{4A}(\boldsymbol{\xi}_{m},\boldsymbol{\xi}_{m}^{\mathsf{c}},\mathbf{w})$. Let $j_{i}^{m}$ and $j_{o}$ denote the inner and outer loop iterations, respectively, for the $m$th AP. Then,
  1. 1.
    At each iteration $(j_{o},j_{i}^{m})$, the optimization variables $\bar{\boldsymbol{\xi}}_{m}^{\;(j_{o},\;j_{i}^{m})}$ and auxiliary variable $\mathbf{w}^{(j_{o},\;j_{i}^{m})}$ satisfy the following sequence of equalities:
    $f_{1A}\big{(}\boldsymbol{\eta}^{(j_{o}+1,\;j_{i}^{m})}\big{)}$ $=f_{2A}\big{(}\boldsymbol{\eta}^{(j_{o}+1,\;j_{i}^{m})}\big{)}=f_{3A}\big{(}\bar{\boldsymbol{\eta}}_{m}^{\;(j_{o}+1,\;j_{i}^{m})}\big{)}$
    $=f_{4A}\big{(}\bar{\boldsymbol{\xi}}_{m}^{\;(j_{o}+1,\;j_{i}^{m})},\mathbf{w}^{(j_{o}+1,\;j_{i}^{m})}\big{)}.$
  2. 2.
    The objective function $f_{4A}\big{(}\bar{\boldsymbol{\xi}}_{m}^{\;(j_{o}+1,\;j_{i}^{m})},\mathbf{w}^{(j_{o}+1,\;j_{i}^{m})}\big{)}$ is monotonically non-decreasing with respect to both the inner and outer loop iterations.
  3. 3.
    The overall objective function $f_{1A}\big{(}\boldsymbol{\eta}^{(j_{o},\;j_{i}^{m})}\big{)}$ converges to a stationary point $\boldsymbol{\eta}^{*}$ as the number of iterations increases.
  4. 4.
    For the original problem $\mathbf{P1}$ with objective function $f_{\text{SE}\;}\big{(}\boldsymbol{\eta},\boldsymbol{\Theta}_{r}\big{)}$, the sequence $f_{\text{SE}\;}\big{(}\boldsymbol{\eta}^{(t)},\boldsymbol{\Theta}_{r}^{(t)}\big{)}$ is monotonically non-decreasing and converges to a sub-optimal point $(\boldsymbol{\eta}^{*},\boldsymbol{\Theta}_{r}^{*})$.
Proof.
The proof directly follows from the application of the equivalent objective conditions [37], equivalent solution properties, first-order optimality conditions [44], and convergence characteristics of the underlying algorithms. ∎
@@ V Simulation Results
In this section, we evaluate the performance of the RIS-assisted cell-free mMIMO NOMA system using extensive simulations. For simulation, we consider a large geographical area of $(1000\times 1000)$ m2, where the APs and users are uniformly distributed. The four vertices of the considered region in terms of $(x,y)$ coordinates are $[-500,500]$ m, $[-500,-500]$ m, $[500,-500]$ m and $[500,500]$ m. The RIS is located at $(x,y)=(0,100)$ m in the considered simulation area. The users are deployed outside a semi-circle of radius $d_{\text{min}}=100$ m from the RIS. We consider the large-scale path-loss parameters according to the three-slope propagation model as $\beta_{mk}=\text{PL}_{mk}+\sigma_{sh}z_{mk}$ (dB[1], where $\text{PL}_{mk}$ is the distance dependent path-loss which depends on the distance between the transmitter and the receiver, $\sigma_{sh}z_{mk}$ represents the shadow fading with standard deviation $\sigma_{sh}$, and $z_{mk}\sim\mathcal{CN}(0,1)$. The shadow fading has a log-normal distribution with a standard deviation ($\sigma_{sh}$) equal to $14$ dB and $4$ dB for the direct and indirect links, respectively. The distance thresholds for the three slopes are $10$ m and $50$ m. The height of the APs, RIS, and users is $15$ m, $30$ m, and $1.65$ m, respectively [13].
@@ V-A Validation of Closed-Form Spectral Efficiency Expression Derived in Theorem-1
Fig. 2(a) illustrates the SE performance for different transmit powers $\rho_{d}$ while serving a total $20$ number of users with two different setups, i.e., $K=10$ clusters with $N=2$ users per cluster and $K=5$ clusters with $N=4$ users per cluster. We set $M=64$ APs, $L=64$ RIS elements, and the coherence interval, $\tau_{c}=100$. We observe from Fig. 2(a) that the derived closed-form SE expression using (18) (labeled as “Closed-form SE (Analytical)”) exactly matches with the SE obtained using (16) (labeled as “Closed-form SE (Simulation)”), which confirms the accuracy of the derivation. We also observe that the derived SE expression using (18) in closely follows the ergodic SE in (13) (labeled as “Ergodic SE (Simulation)”), which validates Theorem 1. To show the advantages of using RIS, we also plot the SE in Fig. 2(a) in the absence of RIS (labeled as “Closed-form SE without RIS”). We clearly observe that for transmitting power $\rho_{d}<0$ dBm, the SE without RIS is almost zero due to the high path loss and shadowing. On the contrary, with the integration of RIS, we get a significant boost in the SE value even with very low transmit power. Therefore, this justifies using the RIS when there is a weak direct link between the user and the AP. We also observe that for higher transmit power, i.e., $\rho_{d}\geq 40$ dBm, the SE obtained with and without RIS are equal. This is because at higher transmit powers, the direct link becomes dominant over the RIS-aided link. As we increase $N$, the SINR at each user decreases due to the increase in intra-cluster interference, which in turn reduces the SE.
V-B Impact of the Number of Users on the Achievable SE
@@ V-C Comparison of Cell-Free NOMA and Cell-Free OMA
V-D Joint Impact of Number of APs and RIS Elements on SE
@@ V-F NOMA-OMA Comparison with the Proposed Algorithm
Figure 5: a) Spectral efficiency versus AP transmit power for different direct path scenarios, b) Spectral efficiency versus AP transmit power for different user distribution scenarios, c) SE versus number of iterations for Algorithm $3$.
V-G Impact of the Number of RIS Elements on the SE with Phase Optimization
@@ V-H Impact of Shadow Fading Coefficient $\sigma_{sh}$ and $d_{\text{min}}$
V-I Convergence of the Proposed Algorithm
To demonstrate the convergence of the proposed algorithm, we now plot in Fig. 5(c) the SE versus the number of iterations required for the Algorithm $3$ to converge, by varying the AP transmit power. For this study, we set $M=64$ APs, $K=10$ clusters, $N=2$ users per cluster and $L=64$ RIS elements. We observe from Fig. 5(c) that Algorithm $3$ requires a small number of iterations (i.e., around $25$ iterations) to converge.
@@ V-K Comparison of RIS Assisted Cell-Free NOMA with SDMA and TDMA
V-L SE versus Users for Different Coherence Interval
@@ VI Conclusion
$\text{IaCI}^{R}_{kn^{\prime}}\stackrel{{\scriptstyle(a)}}{{=}}\sum_{n^{\prime}=n+1}^{N}\sum\limits_{m=1}^{M}{{\eta_{mkn^{\prime}}\rho_{d}}}\delta_{mkn}\gamma_{mk},$ (40)
@@
$\text{ICI}_{k^{\prime}n^{\prime}}=\sum_{k^{\prime}\neq k}^{K}\sum_{n^{\prime}=1}^{N}\sum\limits_{m=1}^{M}{\eta_{mk^{\prime}n^{\prime}}\rho_{d}}{\mathbb{E}\left\{\left|{u}_{mkn}{\hat{z}}_{mk^{\prime}}^{*}\right|^{2}\right\}},$